# ComparisonQA: Evaluating Factuality Robustness of LLMs Through Knowledge Frequency Control and Uncertainty

### v1

@@ 4.2 The Two-Round Measurement
Our ComparisonQA benchmark ensures that the classification and comparison is fair between high-frequency and low-frequency questions. Because we have a shared abstract question for each pair of entities, thus controlling other factors which may affect LLM’s uncertainty and accuracy. Thus, we can rule out all the other factors to purely compare the difference between high-frequency and low-frequency knowledge.
@@ 4.3 Experiments and Analysis
There are very interesting observations from the results: (1) GPT-4o and GPT-4o-mini can not stand the test of robust knowledge measurement especially on low-frequency knowledge: Their performance is still very good on high-frequency questions after the second round but drops a lot on low-frequency ones. The accuracy of GPT-4o on low-frequency questions after the second round is only 25.50, about 61 points lower than the first round. Its performance is even the worst of all the evaluated models, with GPT-4o-mini being the second-to-last. (2) The LLMs’ grasp of low-frequency knowledge is less robust than that of high-frequency knowledge: Performance of LLMs in the second round all drop much more on low-frequency questions than on high-frequency questions. For example, in all of the open source LLMs, the accuracy of Llama-3.1-8B-Instruct in the second round, whose average performance is the best in the first round, dropped up to about 25 points on low-frequency questions while only about 7 points on high-frequency questions. (3) LLMs which can stand the test of robust knowledge measurement must have a very accurate grasp of low-frequency knowledge: LLMs whose average accuracy in the second round is higher than 55 are all those who performed relatively better on the low-frequency questions. This illustrates the importance of the LLMs mastering low-frequency knowledge if they want to be reliable models in the aspect of factual knowledge.
@@ 5 ComparisonQA-Hard
Directly collected questions, especially the high-frequency part, are a bit simple for today’s LLMs. More importantly, they may have low quality and semantic shortcuts. So we introduce a subset called ComparisonQA-Hard for future study, containing only difficult, low-frequency questions that are filtered from ComparisonQA to ensure high quality and avoid semantic shortcuts.
5.1 Simple Low-Quality Question Filtering
The filtering process utilizes both correctness and uncertainty, similar to the last section. Previous benchmarks, like SimpleQA Wei et al. (2024a) and MMLU-Pro Wang et al. (2024a), were collected adversarially based on LLMs’ responses to ensure question difficulty. We enhance this method by also considering LLMs’ uncertainty to achieve the selection of high-quality, shortcut-free questions.
@@ 5.2 Parameters Chosen by Expert Verification
To validate our filtering method, we invite experts to annotate the shortcuts in the 200 randomly sampled questions in the same setting mentioned in §3.3. Results show that 9.4% of the 95.5% correct and high-quality questions are identified as having shortcuts which may decrease difficulty.
@@ 6 Conclusions
In this paper, we first introduce ComparisonQA benchmark to evaluate LLMs’ factual knowledge, with a fully automatic pipeline. This benchmark allows for more controllable and detailed comparisons between high-frequency and low-frequency knowledge of LLMs. Then, we propose a two-round method utilizing correctness and uncertainty to measure LLMs’ knowledge robustness. And we are surprised to find that GPT-4o can not stand such a test, especially on the low-frequency knowledge, and even performs the worst among all LLMs we test. At last, we provide a subset called ComparisonQA-Hard, which contains only difficult low-frequency questions of high quality and no shortcuts for future study. During the question filtering, we also discovered that uncertainty is more helpful in filtering out questions with low quality and shortcuts compared with correctness, which is often used by recent works.
@@
In this paper, we provide a shared abstract question with the entities being the only varying part. This approach ensures that the difference in knowledge frequency between a pair is only related to entity frequency. Future research could investigate methods for measuring the entire knowledge frequency in training data, rather than limiting it to entity frequency. We believe it is a challenging but valuable task. Additionally, our focus is on fixed factual knowledge. But those fast-changing knowledge also deserves attention. Specifically, exploring how knowledge frequency can assist LLMs in acquiring new information is a worthwhile area for further study.
@@ Appendix A Benchmark Statistics
Table 4: The statistics of ComparisonQA benchmark. #Q. Pair refers to the number of question pairs. Hyper. means hypernym. #H. R. and #L. R. refer to the average number of relationships of high frequency entities and low frequency entities respectively.
@@
Appendix B Definition of High and Low Frequency
We randomly sample 1K entities from DBpedia and compute their relationship count. Then we sort these entities in order of their relationship count from highest to lowest. To make contrast significant, entities whose cumulative relationships account for the first $1/3$ of the total number of relationships are defined as the high frequency entities, and those accounting for the last $1/3$ are low frequency entities. Repeating this process 3 times, we get the number of relationships to define high frequency and low frequency entities, which are higher than 185 and lower than 107 separately. Of the 1K randomly sampled entities, 119 entities are classified as high frequent and 621 entities as low frequent.
Appendix C Details in ComparisonQA Construction
In the process of entity pairs extraction, hypernyms that do not have enough entities and low-frequency entities that do not have enough high-frequency entities to pair with are all discarded. Finally, we get 293K entities pairs from 9,261 hypernyms in total. These pairs are then fed into the LLM to generate questions. After LLM’s proofread stage, there are 283K question pairs from 9,166 hypernyms left, which are used to build our ComparisonQA benchmark.
In questions generated by LLMs, the answer for the high frequency entity is $A$, low frequency entity is $B$, and their distractors are $C$ and $D$ respectively. And in the end, we randomly shuffle the 4 options in one question to guarantee the balance of the correct options.
Appendix D Calculation of Uncertainty
Then we compute the uncertainty for each statement. For open-source LLMs, uncertainty refers to their perplexity for generating the correct statement. For proprietary LLMs, we allow them to generate their uncertainty scores, as their perplexity is not accessible. The threshold between high and low uncertainty for each model is determined by its own average uncertainty on the benchmark respectively.
@@ Appendix E Experiment Details
Table 7: Performance of various LLMs on the testing set of ComparisonQA. Unc., Acc, and Ma-F1, denote Uncertainty, Accuracy, and Macro F1-score. And the Difference column shows how scores change from high frequency questions to low frequency questions. The best performances within each method are underlined and the best among all methods are bold-faced. And for the Difference column, We underline the largest difference within each method and bold the one among all methods.
For the main evaluations on ComparisonQA, we categorize the evaluation of different models into three types: (1) Open Source LLM Zero-Shot: We first evaluate Llama3, Llama3.1, Llama3.2 Touvron et al. (2023); Dubey et al. (2024), Gemma2 Mesnard et al. (2024); Riviere et al. (2024), Phi3.5 Abdin et al. (2024), Falcon, Falcon2 Malartic et al. (2024), Mistral Jiang et al. (2023), and their instruction versions accordingly in a zero-shot manner Qin et al. (2023). (2) Open Source LLM Few-Shot: Then we evaluate the above models in a few-shot manner Brown et al. (2020). Since our benchmark is in the form of four options multiple-choice questions, the shot number is set to four. (3) Proprietary LLM API: Finally, we evaluate the performance of GPT-4o OpenAI (2023, 2024b) and GPT-4o-mini OpenAI (2024a), using zero-shot, few-shot, and Chain-of-Thought (CoT; Wei et al., 2022). All the evaluation results are reported in Table 7.
And for the question filtering of ComparisonQA-Hard, the six open-source LLMs we use are Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, gemma-2-9b, Phi-3.5-mini-instruct, Falcon-11B, and Mistral-7B-Instruct-v0.3.
All the open-source models are run on 4 NVIDIA A6000 (40G) GPUs with BF32. And for proprietary LLM, we access them via OpenAI API  [^2] https://platform.openai.com/docs/api-reference. The different kinds of prompts we use are shown in Table 6.

### v2

@@ 4.2 The Two-Round Measurement
Although the second round of evaluation is stricter, our ComparisonQA benchmark ensures fair classification and comparison between high-frequency and low-frequency questions under the same setting. Because we have a shared abstract question for each pair of entities, controlling other factors which may affect LLM’s uncertainty and accuracy. Thus, we can rule out all the other factors to purely compare the difference between high-frequency and low-frequency knowledge.
@@ 4.3 Experiments and Analysis
There are very interesting observations from the results: (1) GPT-4o and GPT-4o-mini can not stand the test of robust knowledge measurement especially on low-frequency knowledge: Their performance is still very good on high-frequency questions after the second round but drops a lot on low-frequency ones. The accuracy of GPT-4o-mini on low-frequency questions after the second round is only 38.25, about 35 points lower than the first round. Their average performance even drop the most among all the evaluated models. (2) The LLMs’ grasp of low-frequency knowledge is less robust than that of high-frequency knowledge: Performance of LLMs in the second round all drop much more on low-frequency questions than on high-frequency questions. For example, in all of the open source LLMs, the accuracy of Llama-3.1-8B-Instruct in the second round, whose average performance is the best in the first round, dropped up to about 25 points on low-frequency questions while only about 7 points on high-frequency questions. (3) LLMs which can stand the test of robust knowledge measurement must have a very accurate grasp of low-frequency knowledge: LLMs whose average accuracy in the second round is higher than 55 are all those who performed relatively better on the low-frequency questions. This illustrates the importance of the LLMs mastering low-frequency knowledge if they want to be reliable models in the aspect of factual knowledge.
@@ 5 ComparisonQA-Hard
Directly collected questions, especially the high-frequency part, are a bit simple for today’s LLMs. More importantly, they may have low quality and semantic shortcuts. So we introduce a new method to select a subset called ComparisonQA-Hard, containing only difficult, low-frequency questions that have high quality and no semantic shortcuts.
5.1 Hard High-Quality Question Filtering
The filtering method utilizes both correctness and uncertainty, similar to the last section. Previous benchmarks, like SimpleQA Wei et al. (2024a) and MMLU-Pro Wang et al. (2024a), were collected adversarially based on LLMs’ responses to ensure question difficulty. We enhance this method by also considering LLMs’ uncertainty to achieve the selection of high-quality, shortcut-free questions.
@@ 5.2 Parameters Chosen by Expert Verification
To validate our filtering method, we invite experts to annotate the shortcuts in the 200 randomly sampled questions in the same setting mentioned in §3.3. The instruction is the same as the standard (2) given to LLMs above. Results show that 9.4% of the 95.5% correct and high quality questions are identified as having semantic shortcuts.
@@ 6 Conclusions
In this paper, we first introduce ComparisonQA benchmark to evaluate LLMs’ factual knowledge, with a fully automatic pipeline. This benchmark allows for more controllable and detailed comparisons between high-frequency and low-frequency knowledge of LLMs. Then, we propose a two-round method utilizing correctness and uncertainty to measure LLMs’ knowledge robustness. And we are surprised to find that even powerful LLMs like GPT-4o can not stand such a test, especially on the low-frequency knowledge. At last, we discover that uncertainty is more effective in filtering out questions with low quality and shortcuts compared with correctness, which is often used by recent works. Based on this method, we provide a subset called ComparisonQA-Hard, which contains only difficult low-frequency questions of high quality and no shortcuts for future study.
@@
In this paper, we provide a shared abstract question with the entities being the only varying part. However, due to the sharing of such abstract questions, it is difficult to contain more entities with different frequencies in the same question because many entities lack sufficient shared features to generate common questions.
Our approach ensures that the difference between a pair is only the entity. Future research could investigate methods for measuring the entire knowledge frequency required to solve the questions, rather than limiting it to entity frequency. We believe it is a challenging but valuable task.
@@
Acknowledgments
The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong,SAR, China.
@@ Appendix A Benchmark Statistics
Table 4: The statistics of ComparisonQA benchmark. #Q. Pair refers to the number of question pairs. Hyper. means hypernym. #H. R. and #L. R. refer to the average number of relationships of high-frequency entities and low-frequency entities respectively.
@@
For benchmark quality, we further compute the correctness of the annotated date in §3.3 coming from the train, validation, and test set separately. The results presented in Table 6 show the high correctness in all splits of our benchmark. Besides, our labeling is very strict. If either of the questions in one pair is wrong, we will judge the whole question pair incorrect. Since each pair contains two questions, actually only about 2.25% of the total questions are incorrect.
Split Correctness Agreement
Train 95.68 88.27
Valid 93.75 87.50
Test 95.45 90.91
Task Prompt
Question Generation You will be given two entities belonging to the same hypernym. Generate a shared multiple choice question for both entities based on their descriptions according to the following 5 steps. Requirements: First, the questions should have one and only one correct answer for both entities. Second, the correct answer cannot be simply guessed by the names of the entities or the way the questions are asked. For example, the question: "What was the primary operational location of Sydney O-Class Tram? A. Oslo, Norway B. Sydney, Australia C. Stockholm, Sweden D. Melbourne, Australia" is not allowed since only the correct answer contains "Sydney" which is also in the entity’s name. Third, the four options should have roughly the same length. Step 1, generate the shared question containing "[entity_name]" which can be replaced by the two entity names and then have different answers accordingly. Step 2, use roughly the same amount of words to answer the questions for both entities separately. Step 3, generate a misleading distractor for both entities separately. The distractors should have roughly the same length with the correct answers. Step 4, form the final multiple choice question using the above four answer candidates, and make sure the answer for Entity1 is ’A’ , the answer for Entity2 is ’B’, and their misleading answer candidates are ’C’ and ’D’. Step 5, check whether the final question and answer candidates meet the above requirements. If yes, then output **SCUUEED**, otherwise output **FAIL**.
Follow these examples: ……(Examples written by experts)
Hypernym: [Hypernym] Entity1: [Entity1] Entity1 Description: [Entity1 Description] Entity2: [Entity2] Entity2 Description: [Entity2 Description]
Table 7: The prompt used to generate questions in ComparisonQA. Placeholders [Hypernym], [Entity1], [Entity1 Description], [Entity2], [Entity2 Description] will be replaced with real hypernym, high-frequency entity, low-frequency entity, and their descriptions accordingly.
Appendix B Definition of High and Low Frequency
First, we randomly sample 1K entities from DBpedia and compute their relationship count separately. This process is to study the distribution of entities’ relationship counts in DBpedia and find boundaries for high-frequency and low-frequency entities. Then we sort these entities in order of their relationship count from highest to lowest. High-frequency entities possess cumulative relationship count of up to 1/3 of all entities, while low frequency entities range from 2/3 to 1. We exclude those between 1/3 and 2/3 to make comparison clear.
Repeating this process 3 times, we get the number of relationships to distinguish high-frequency and low-frequency entities, which are higher than 185 and lower than 107 respectively. (Of the 1K randomly sampled entities, 119 entities are identified as high frequent and 621 entities as low frequent.) These two numbers are then used to classify all the high-frequency and low-frequency entities.
It is necessary to randomly select 1K entities first to calculate the two boundaries, since the time to compute the cumulative relationship count for all the sorted entities in DBpedia is unaffordable.
Appendix C Details in ComparisonQA Construction
In the process of entity pairs extraction, both hypernyms that do not have enough entities and low-frequency entities that do not have enough high-frequency entities to pair with are all discarded. Finally, we get 293K entities pairs from 9,261 hypernyms in total.
These pairs are then fed into the LLM to generate questions. Each pair has a shared abstract question respectively. Since a hypernym can have several entity pairs, the final number of abstract questions, which is equal to the number of pairs, is more than that of hypernyms. After LLM’s proofread stage, there are 283K abstract questions left, which are used to build our ComparisonQA benchmark. Entities belonging to these questions come from 9,166 different hypernyms.
In the process of Abstract Question Generation, the prompt we use to generate questions is shown in Table 7. In questions generated by LLMs, the answer for the high-frequency entity is $A$, low-frequency entity is $B$, and their distractors are $C$ and $D$ respectively. And in the end, we randomly shuffle the 4 options in one question to guarantee the balance of the correct options.
Task Prompt
Uncertainty Question: [question]
Generation Answer: **[option].**
Uncertainty: **[uncertainty percentage]**
Answer the following multiple choice question. Select only one correct answer from the choices and give your uncertainty score, following the above format.
[Question]
A. [OptionA]. B. [OptionB].
C. [OptionC]. D. [OptionD].
Table 8: The prompt used to generate uncertainty score for proprietary LLMs. Placeholders [Hypernym], [Entity1], [Entity1 Description], [Entity2], [Entity2 Description] will be replaced with real hypernym, high-frequency entity, low-frequency entity, and their descriptions accordingly.
Appendix D Calculation of Uncertainty
Then we compute the uncertainty for each statement. For open-source LLMs, uncertainty refers to their perplexity for generating the correct statement. For proprietary LLMs, we allow them to generate their uncertainty scores, with prompt shown in Table 8. Then, the threshold between high and low uncertainty for each model is determined by its own average uncertainty on the testing set of each benchmark respectively.
There are several reasons for choosing different uncertainties for each setting. On one hand, perplexity-based uncertainty is unavailable for proprietary LLMs. On the other hand, open source LLMs often fail to understand the instructions and can not generate uncertainty scores. We have conducted experiments on verbalized uncertainty of open source LLMs. However, experiments show that Llama-3-8B-Instruct has a 18.5% chance of not generating an uncertainty score, which makes the results unreliable and will also affect further robustness evaluations. In addition, Vashurin et al. (2024) mentioned that, for multiple choice QA, information-based methods such as perplexity are substantially superior to quantifying model uncertainty.
@@ Appendix E Experiment Details
Table 10: Performance of various LLMs on the testing set of ComparisonQA. Unc., Acc, and Ma-F1, denote Uncertainty, Accuracy, and Macro F1-score. And the Difference column shows how scores change from high-frequency questions to low-frequency questions. The best performances within each method are underlined and the best among all methods are bold-faced. And for the Difference column, We underline the largest difference within each method and bold the one among all methods.
For the main evaluations on ComparisonQA, we categorize the evaluation of different models into three types: (1) Open Source LLM Zero-Shot: We first evaluate Llama3, Llama3.1, Llama3.2 Touvron et al. (2023); Dubey et al. (2024), Gemma2 Mesnard et al. (2024); Riviere et al. (2024), Phi3.5 Abdin et al. (2024), Falcon, Falcon2 Malartic et al. (2024), Mistral Jiang et al. (2023), and their instruction versions accordingly in a zero-shot manner Qin et al. (2023). (2) Open Source LLM Few-Shot: Then we evaluate the above models in a few-shot manner Brown et al. (2020). Since our benchmark is in the form of four options multiple-choice questions, the shot number is set to four to minimize bias, where each of the four examples corresponds to a different correct answer in (a, b, c, d). (3) Proprietary LLM API: Finally, we evaluate the performance of GPT-4o OpenAI (2023, 2024b) and GPT-4o-mini OpenAI (2024a), using zero-shot, few-shot, and Chain-of-Thought (CoT; Wei et al., 2022).
All the open-source models are run on 4 NVIDIA A6000 (40G) GPUs with BF32. And for proprietary LLM, we access them via OpenAI API  [^2] https://platform.openai.com/docs/api-reference. The different kinds of prompts we use are shown in Table 9. And all the evaluation results are reported in Table 10.
For the question filtering of ComparisonQA-Hard, the six open-source LLMs we use are Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, gemma-2-9b, Phi-3.5-mini-instruct, Falcon-11B, and Mistral-7B-Instruct-v0.3.